Eckel recently set up a TDA to save for his retirement. He arranged to have $130 taken out of each of his biweekly checks; it will earn 9 7/8% interest. He just had his twenty-ninth birthday, and his ordinary annuity comes to term when he is 65. Find the following. (Round your answers to the nearest cent.)

(a) The future value of the account
$  

(b) Eckel's total contribution to the account
$  

(c) The total interest
$
Answer #1

1.
The future value of the account=130/(9.875%/26)*((1+9.875%/26)^(26*36)-1)=1155292.6338

2.
Eckel's total contribution to the account=130*26*36=121680

3.
The total interest=1155292.6338-121680=1033612.6338